Can we expect a precipitate of lead (II) iodide, PbI2, to form when 20.00 mL of 0.0050 mol/L 

     aqueous calcium iodide, CaI2, solution is added to 80.0 mL of 0.0010 mol/L aqueous solution of  lead(II) nitrate, Pb(NO3)2?  The Ksp for lead (II) iodide is 1.4X10-8. 


1
Expert's answer
2021-01-15T06:30:43-0500

The ionic equation for Cal, is as follows:

"CaI_{(aq)}\u2192 2I^{-}_{(aq)} + Ca^{2+}_{(aq)}"


The ionic equation for Pb(NO3)2, is as follows:

Pb(NO3)2(aq) → Pb2+(aq) + 2NO-3 (aq)



The ionic equation for solubility is as follows:

Pb(NO3)2(aq) + Cal(aq) → PbI(s)+ Ca(NO3)2(aq)


The reaction quotient Q is:

Q = [Pb2+][I-]²



The initial concentration of Pb2+ is as follows:

Initial concentration of [Pb2+] = "\dfrac{Molarity×Volume}{Total \ volume}" = "\dfrac{0.0010M × 80.0 mL}{80.0mL+20.0 mL}" = 1.0×10-3M



The reaction quotient Q will be as follows:

Q = [Pb2+][I-]²

= [8x10-4] 1.0x10-3]2

= 8.0 x10-10M


The solubility product of Pbl2 is 1.4 x10-8 and the reaction quotient is 8.0x10-10 is less than the Ksp

So, Ksp > Q therefore precipitation does not occur